OpenStudy (anonymous):

4sqrt(16)-16sqrt(8)+sqrt(4)=?

OpenStudy (anonymous):

show how to solve please :)

OpenStudy (hoblos):

\[4\sqrt{16} - 16\sqrt{8} +\sqrt{4} = 4*4 -16*2\sqrt{2} +2= 18-32\sqrt{2}\]

OpenStudy (anonymous):

you are awesome :) thank you

OpenStudy (hoblos):

my pleasure XD